angle a is complementary to angle c. if angle a is equal to 2x-10 and angle c is equal to 4x+40, find the measurement angle of a, the measurement angle of c, and the measurement angle of b
Answer by RichardMathsTutor(3)

If angle a is complimentary to angle c then angle a and angle c add up to 90 degrees.
So angle a + angle c = 90
If angle a = 2x-10 and angle c = 4x+40
Then 2x-10+4x+40 = 90
Simplifying LHS gives 6x+30 = 90, and thus 6x = 60 or x = 10
Substituting this value of x into 2x-10 and 4x+40 gives:
angle a = 2(10)-10 = 10, angle c = 4(10)+40 = 80
To check; angle a + angle c = 10+80 = 90, this is complimentary.
More information is needed regarding angle b.